9. Calculate ΔG0 at 820C for reactions in which

(a) ΔH0= 293 kJ; ΔS0 = -695 J/K

(b) ΔH0= -1137 kJ; ΔS0= 0.496 kJ/K
Answer #1

G0 = H0 -TS0

We use this formula for both the below questions

T = 820C = (82+273)K = 355K

(a) H0 = 293kJ = 293000J [Since 1kJ = 1000J]

G0 = H0 -TS0

G0 = 293000J - (355K x -695J/K)

G0 = 293000J + 246725J = 539725J

(b) H0 = -1137kJ

G0 = H0 -TS0

G0 = -1137kJ - (355K x 0.496kJ/K)

G0 = -1137kJ - 176.08kJ

G0 = 1313.08kJ = 131080J